OpenStudy (anonymous):

the cost of manufacturing clocks is given by c=36(n+81)1/2, where c is the cost in dollars and is the number produced. what is the cost when no clocks is produced?

OpenStudy (anonymous):

\[36(n+81)^{1/2}=36\sqrt{n+81}\] \[36\sqrt{0+81}\]

OpenStudy (anonymous):

You'd enter 0 for n, leaving you with cost= 36x sqrt(81^1). 81 to 1/2 power is the squAre root of 81 to the first power. Anything to the first power is itself. The square root of 81 is 9. 9 x 36 = 224. 224 is the cost with nothing produced.

OpenStudy (anonymous):